# Intent-Aware Dialogue Generation and Multi-Task Contrastive Learning for Multi-Turn Intent Classification

### v1

@@ Abstract.
Generating large-scale, domain-specific, multilingual multi-turn dialogue datasets remains a significant hurdle for training effective Multi-Turn Intent Classification models in chatbot systems. In this paper, we introduce Chain-of-Intent, a novel mechanism that combines Hidden Markov Models with Large Language Models (LLMs) to generate contextually aware, intent-driven conversations through self-play. By extracting domain-specific knowledge from e-commerce chat logs, we estimate conversation turns and intent transitions, which guide the generation of coherent dialogues. Leveraging LLMs to enhance emission probabilities, our approach produces natural and contextually consistent questions and answers. We also propose MINT-CL, a framework for multi-turn intent classification using multi-task contrastive learning, improving classification accuracy without the need for extensive annotated data. Evaluations show that our methods outperform baselines in dialogue quality and intent classification accuracy, especially in multilingual settings, while significantly reducing data generation efforts. Furthermore, we release MINT-E, a multilingual, intent-aware multi-turn e-commerce dialogue corpus to support future research in this area.
Multi-turn Intent Classification, Large Language Model, Multi-task Contrastive Learning
copyright: ccjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Language modelsccs: Computing methodologies Intelligent agentsccs: Information systems Question answering
@@
**footnotetext: Equal Contributions.$\dagger$\$$\dagger$\$footnotetext: Corresponding Author: bin.fu@shopee.com
Figure 2. Example of intent-aware multi-turn dialogue and the challenges of tagging single-turn cross-market intents to create multi-turn intent database. The intents (e.g. Chitchat - Complaint) are in the format of ¡intent type¿-¡intent name¿.
@@ 1. Introduction
Chatbots are essential on international e-commerce platforms, offering 24/7 customer service. Users typically engage with these chatbots through multi-turn conversations to achieve their goals. Therefore, it is crucial for the chatbot to accurately identify the user’s intent at each turn using a Multi-Turn Intent Classification (MTIC) model. To improve accuracy, MTIC models should consider not just the current message but also previous interactions and intents. Ignoring the context from earlier turns often leads to mistakes, giving irrelevant answers and making users dissatisfied (Xu and Sarikaya, 2014).
@@
Motivation and Challenges. Similar to many deep learning approaches (Soudani et al., 2024), MTIC models heavily rely on large amounts of domain-specific multi-turn intent-utterance pairs to boost their performance (Ni et al., 2023). However, gathering such datasets efficiently, especially across different languages, is challenging. While some studies on dialogue understanding in multi-turn intent classification (Wu et al., 2021; Qu et al., 2019; Ren and Xue, 2020) assume there is plenty of multi-turn training data available, that is often not the case in reality. Recent work (Zhao et al., 2023) shows that using few-shot prompting with ChatGPT does not perform well for intent classification, and it does not match up to supervised methods.
@@
Figure 2 illustrates the difficulties in annotating multi-turn dialogues. For MTIC in information-seeking dialogues (Askari et al., 2024a), manual annotation of intents is time-consuming and resource-intensive, which leads to a lack of training data. Unlike popular intent classification datasets such as MSdialog-Intent (Qu et al., 2018), which has only 12 unique intents, industrial chatbots handle hundreds of intents to meet specific user needs in each market. This vast number of intents makes classification tasks and data annotation more complex. With so many options, annotators can easily make errors and need more time to make the right decisions, making the process expensive and slow. For instance, Qu et al. (2019) reported that labelling intents for 2,199 dialogues cost \$1,700. Manually annotating large-scale, multi-turn datasets in multiple languages or constantly updating the model for new user intents is impractical. Lack of sufficient training data often leads to poorer performance, highlighting the need for a better way to address this challenge.
@@
Most research integrates the dialogue acts (DA) for domain-specific task-oriented dialogue (TOD) generation (Wang et al., 2020b). However, dialogue acts are more general, like statement/request/yes-answer, and therefore unusable for domain-specific generation. Recent works also adopt specific DA in the e-commerce domain for only TOD intents (Chen et al., 2019). However, there are FAQ, chit-chat, and TOD intents in multi-turn sessions shown in Figure  2, only considering that TOD intents are limited for dialogue generation in real-world scenarios.
@@
To tackle these issues, this paper contributes three solutions from different aspects, as follows:
@@ 2.1. Single-turn
In a closed domain question answering (QA) system, we have a set of predetermined intent classes =$\mathcal{I}=\{I_{i}\}_{i=1}^{k}$, ranging from informational queries to action-oriented requests. Here, we assume that a dataset of single-turn question-intent pairs $\mathcal{D}={(x_{i},y_{i})}$ can be easily accessible, where $x_{i}$ denotes the question and $y_{i}\in\mathcal{I}$ is its labelled intent. This dataset is described as single-turn because all $x_{i}$ on their own are standalone complete utterances and are semantically clear for intent inference.
@@ 2.2. Multi-turn
Figure 3. Detailed illustration of using LLM-enhanced HMM to sample a chain of intents and generate intent-aware multi-turn dialogues
@@ 3. Intent-aware Dialogue Generation
Figure 3 shows the detail workflow of intent-aware dialogue generation, which comprises two main stages: domain knowledge extraction and question generation.
@@ 3.1. Domain Knowledge Extraction
As LLMs do not possess domain-specific knowledge out-of-the-box, this step is imperative to extract the knowledge to be injected into the conversation generation process, bridging the gap for business usage. There are two domain-specific knowledge to account for when generating a conversation, namely the number of turns in each conversation and the intent transition from intents of historical questions. These two combined will guide the generation of each conversation.
@@ 3.1.2. Intent Transition Matrix
Other than the number of turns, we also need to know the sensible sequence of intents that users might have in a conversation. Just like the number of turns, the intent transition information is also gathered from the chat logs. One limitation here is that the intent from chat logs is inferred by an intent recognition model trained on the single-turn data $\mathcal{D}$. Each $q_{t}$ is inferred separately without considering $h_{t}$ to obtain $I_{t}$ for the $t$-th turn. That said, by the Law of Large Numbers, if we gather the statistics from a reasonable large number of chat logs, the intent transition matrix obtained will provide a good estimation to true intents. Hereby, the initial intent distribution $P(I_{1})$ is specified in a row matrix, while the intent transition $P(I_{t}|I_{t-1})$ is defined in another $|\mathcal{I}|\times|\mathcal{I}|$ matrix. Note that the $P(I_{t}|I_{t-1})$ here is time invariant, the probability of $P(I_{j}|I_{i})$ will be the same regardless of $t$.
@@ 3.2.1. Intent Sampling with HMM
Using these distributions, we sample the number of turns $T$ from $P(T)$, the initial intent $I_{1}$ from $P_{\text{init}}$, and subsequent intents $I_{t}$ from $P_{\text{trans}}$ based on $I_{t-1}$.
@@
(1) $P(S_{1:T},Y_{1:T})=P(S_{1})P(Y_{1}|S_{1})\prod_{t=2}^{T}P(S_{t}|S_{t-1})P(Y_{t}|S_{t})$
In our case, $S_{t}$ corresponds to the intent $I_{t}$, and $Y_{t}$ corresponds to the user’s utterance (question) at turn $t$. The observation model $P(Y_{t}|S_{t})$ is modeled by uniformly sampling a question $X_{t}$ from the set of questions in the single-turn dataset $\mathcal{D}$ that are labeled with intent $I_{t}$:
@@ 3.2.3. Enhancing HMMs with LLMs
To overcome these limitations, we integrate LLMs into the HMM framework, allowing us to generate dialogues that are coherent and context-aware. Specifically, we leverage the LLM’s in-context learning ability to generate each user utterance $q_{t}$ and the corresponding chatbot response $a_{t}$, conditioning on the conversation history up to turn $t-1$.
@@
Formally, we generalize Equation 2 to:
@@
×$\times\prod_{t=2}^{T}P(I_{t}|I_{t-1})P(X_{t}|I_{t})P(a_{t-1}|q_{1:t-1})$
(3) ×$\times P(q_{t}|I_{t},X_{t},q_{1:t-1},a_{1:t-1})$
@@ 3.3. MINT-E
As ID is the main market with the most traffic online, we observe the top 15 intents based on the question-level ratio in ID generated dataset. The top user intent is to check delivery status, which makes up to 13.8% of the questions. We also observed a long-tailed distribution with the top 10 intents taking up more than 50% of the questions. The detail distribution is included in Appendix.
@@
Table 1. Statistics of UDC, MSDialog, and MINT-E.
@@ 4.3.3. Combined Training Objective
Here, $\lambda$ is a hyperparameter that balances the contribution of the contrastive loss. Through manual tuning, we found that setting $\lambda=0.3$ yields the best results.
@@ 5.2. Settings for Dialogue Quality Evaluation
We evaluated the quality of the synthetic dialogues using the GPT-4 model. The assessment focused on language fluency, logical flow of topics, and the coherence of questions within each session. Conversations were rated on a scale from 1 to 10, where 1 indicated poor quality and 10 represented excellent quality. The overall score was calculated by averaging the ratings across all evaluated conversations. The prompt for this evaluation can be found in Appendix A.4.
@@ 5.4.1. Quality of Generated Dialogues
Table 3 shows the quality ratings for the generated dialogues across different markets. Our generated dataset, MINT-E, consistently outperformed the baselines in every market. Interestingly, our generated dialogues sometimes received higher average ratings than the golden set. As users often provide brief or out-of-context queries, whereas the language model generates more detailed and contextually relevant questions.
@@
Table 3. Quality ratings of generated dialogues across different methods and markets. The best performance is highlighted in bold.
@@
Table 4. Impact of different training configurations on Multi-Turn Intent Classification accuracy. Best performance (averages of all markets and English (EN) market only) is highlighted in bold.
@@ 5.4.2. Downstream Task Performance
The impact of different training configurations on the MTIC task is presented in Table 4. Incorporating multi-turn data generally improved performance compared to using only single-turn data. The multi-task learning strategies improved average accuracy by 0.5% across 5 markets, particularly in English-speaking markets.
@@ 5.4.3. Discussion and Limitations
Our experiments show that integrating domain knowledge with LLM, as done in MINT-CL, noticeably enhances dialogue quality and improves performance on downstream tasks. The results suggest that leveraging multi-turn data and multi-task learning is beneficial, especially in markets where the language model performs well.
@@
Table 5. Single-turn (ST) and Multi-turn (MT) result on low resources languages
@@
Table 6. Statistics of MINT-E compared to golden set sampled from true online chat logs.
@@ 7. Related Work
Multi-turn Dialogue Generation (MTDG) emerge from both academics and industry due to its relevance to real-world applications such as chatbots and information-seeking systems (Liu and Fu, 2024). However, they face challenges such as the data sparsity (Zhang et al., 2018c, b; Li et al., 2017). Moreover, MTDG requires more complex information and constraints (Zhang et al., 2020; Chen et al., 2018a; Zhang et al., 2018a), posing additional challenges. In general, dialogue generation are categorised in open-domain generation and task-oriented generation (Lv et al., 2023).
@@
Open-domain generation is a context-sensitive process that spans multiple turns, with the model learning to generate appropriate yet open-ended responses based on previous utterances (i.e., the context), which can be categorized into preLLM and postLLM (Soudani et al., 2023). PreLLM methods include the HRED (Serban et al., 2015) series of models. This hierarchical encoder-decoder framework effectively captures contextual information and incorporates latent variables as intermediate states to generate diverse responses, such as variable HRED (Serban et al., 2016b), MrRNN (Serban et al., 2016a) and DialoGPS(Lv et al., 2023). Other related research adopts deep-learning models (Santhanam and Shaikh, 2019) enhanced by internal knowledge to ensure on-topic response generation(Yu et al., 2022; Madotto et al., 2018; Wang et al., 2020a). Popular architecture includes recurrent neural network(RNN) (Li et al., 2015), memory network (Chen et al., 2018b) and transformer (Lin et al., 2020; Olabiyi et al., 2023). Those works introduce topic words (Xing et al., 2017), emotion (Song et al., 2019; Zhou et al., 2018), and persona (Zhang, 2018; Zhou et al., 2018) into the generation process to help deliver better dialogue responses. Most of the research still adopts supervised models, which require much training data for better performance. The advent of LLM has created new possibilities for automated text generation. PostLLM methods leverage powerful LLM to utilize prompt or instruction to generate instruction-response pairs automatically. Due to the excellent generation ability of LLM, most research on generation tasks focuses on developing high-quality instruction for multi-turn generation (Wang et al., 2022; Ding et al., 2023) and enhancing the multi-turn instruction following ability with human feedback (Sun et al., 2023; Ouyang et al., 2022). However, those open-domain generation methods didn’t explicitly adopt the domain-specific intent into dialogue generation, so their work is hardly used in other downstream NLP tasks, such as intent or emotion classification tasks.
@@
Task-oriented generation is goal-driven and limited to specific domains for task-oriented dialogue(TOD) (Soudani et al., 2023). It has been widely deployed in many industries as it delivers more efficient customer support (Moghe et al., 2022; Deng et al., 2022). The popular approach centres around user simulation (Lin et al., 2021; Tseng et al., 2021), where two agents interact to process the dialogue history and create a belief state for each side of the conversation, called self-play. They adopt HMM (Cuayáhuitl et al., 2005), partially observable Markov decision processes (POMDP), as a user model for spoken dialogue generation. Recent work also adopts LLM as a user simulator (Kong et al., 2024) with reinforcement learning (Min et al., 2023) and in-context learning framework (Terragni et al., 2023; Askari et al., 2024b). This work considers the limited number of dialogue acts, which can’t include all types of intents within the chatbot.
@@
Multi-Turn Intent Classification focuses on understanding user intents within the context of ongoing dialogues. Larson et al. (2019) highlighted the challenge of distinguishing out-of-scope queries, emphasizing the need for robust intent classification models. To enhance interpretability in intent identification, Yang et al. (2020) utilized reinforcement learning over knowledge graphs, demonstrating improved performance in multi-turn settings. Tran et al. (2022) proposed an adaptive global-local context fusion mechanism, effectively integrating dialogue context for enhanced intent classification and slot filling. To address the complexity of multi-turn conversational contexts in the e-commerce domain, Liu et al. (2024) developed LARA, a Linguistic-Adaptive Retrieval-Augmentation framework across multiple languages. Furthermore, Fan et al. (2022) developed a system leveraging pre-trained language models for fine-grained query understanding, aligning with our focus on intent-aware dialogues in specific domains. Additionally, Hao et al. (2023) introduced a method employing reinforcement learning to construct an intent graph for identification, highlighting the importance of capturing intent transitions in dialogues.

### v2

# From Intents to Conversations: Generating Intent-Driven Dialogues with Contrastive Learning for Multi-Turn Classification

@@ Abstract.
In conversational AI systems, a critical challenge in training effective multi-turn intent classification models lies in the generation of large-scale, domain-specific, multilingual dialogue datasets. In this paper, we introduce Chain-of-Intent, a novel framework that integrates Hidden Markov Models (HMMs) with Large Language Models (LLMs) to generate intent-driven, context-aware dialogues through self-play. Our method first extracts domain-specific intent transition patterns from real-world e-commerce chat logs, which guide the modeling of turn-level dynamics and intent sequences. LLMs are then employed to parameterize the emission probabilities of HMMs, enabling the generation of natural, coherent utterances aligned with predicted intents and dialogue context. We further propose MINT-CL, a multi-task contrastive learning framework for multi-turn intent classification, which improves performance while reducing dependence on large-scale annotated datasets. Empirical results demonstrate that our approach outperforms competitive baselines in both dialogue generation quality and classification accuracy, particularly in multilingual settings. To facilitate future research, we release MINT-E, a comprehensive, multilingual, intent-aware multi-turn dialogue corpus derived from the e-commerce domain [^1] The reproduced source code and dataset are available at https://github.com/junhua/chain-of-intent..
Multi-turn Intent Classification, Large Language Model, Multi-task Learning, Contrastive Learning
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 34th ACM International Conference on Information and Knowledge Management; November 10–14, 2025; Seoul, Republic of Koreabooktitle: Proceedings of the 34th ACM International Conference on Information and Knowledge Management (CIKM ’25), November 10–14, 2025, Seoul, Republic of Koreadoi: 10.1145/3746252.3761117isbn: 979-8-4007-2040-6/2025/11ccs: Information systems Language modelsccs: Computing methodologies Intelligent agentsccs: Information systems Question answering
@@
Figure 2. Example of intent-aware multi-turn dialogue and the challenges of annotating single-turn cross-market intents to create a multi-turn intent classification dataset. The intents (e.g., Chitchat - Complaint) are in the format of ¡intent type¿-¡intent name¿.
@@ 1. Introduction
Chatbots are essential on international e-commerce platforms, offering 24/7 customer service. Users typically interact with these chatbots through multi-turn conversations to achieve their goals. Therefore, it is crucial for the chatbot to accurately identify the user’s intent at each turn using a Multi-Turn Intent Classification (MTIC) model. To improve accuracy, MTIC models should consider not just the current message but also previous interactions and intents. Ignoring the context from earlier turns often leads to mistakes, giving irrelevant answers and making users dissatisfied (Xu and Sarikaya, 2014).
@@
Motivation and Challenges. Like many deep learning approaches (Soudani et al., 2024), MTIC models heavily rely on large amounts of domain-specific multi-turn intent-utterance pairs to boost their performance (Ni et al., 2023). However, collecting such datasets efficiently, especially across different languages, is challenging. While some studies on dialogue understanding in multi-turn intent classification (Wu et al., 2021; Qu et al., 2019; Ren and Xue, 2020) assume there is plenty of multi-turn training data available, that is often not the case in reality. Recent work (Zhao et al., 2023) shows that using few-shot prompting with ChatGPT does not perform well for intent classification, and it does not match up to supervised methods.
@@
Figure 2 illustrates the challenges in annotating multi-turn dialogues. For MTIC in information-seeking dialogues (Askari et al., 2024), manual annotation of intents is time-consuming and resource-intensive, which leads to a lack of training data. Unlike popular intent classification datasets such as MSdialog-Intent (Qu et al., 2018), which has only 12 unique intents, industrial chatbots handle hundreds of intents to meet specific user needs in each market. This vast number of intents significantly increases the complexity of both classification tasks and data annotation. With so many options, annotators can easily make errors and need more time to make the right decisions, making the process expensive and slow. For instance, Qu et al. (2019) reported that labelling intents for 2,199 dialogues cost \$1,700. Manually annotating large-scale, multi-turn datasets in multiple languages or constantly updating the model for new user intents is impractical. Lack of sufficient training data often leads to poorer performance, highlighting the need for a better way to address this challenge.
@@
Most research integrates the dialogue acts (DA) for domain-specific task-oriented dialogue (TOD) generation (Wang et al., 2020b). However, dialogue acts are more general, like statement/request/yes-answer, and therefore unusable for domain-specific generation. Recent works also adopt specific DA in the e-commerce domain for only TOD intents (Chen et al., 2019). However, there are FAQ, chit-chat, and TOD intents in multi-turn sessions shown in Figure  2, so considering only TOD intents is limited for dialogue generation in real-world scenarios.
@@
To address these issues, this paper contributes three solutions from different aspects, as follows:
@@ 2.1. Single-turn
In a closed domain question answering (QA) system, we have a set of predetermined intent classes $\mathcal{I}=\{I_{i}\}_{i=1}^{k}$, ranging from informational queries to action-oriented requests. Here, we assume that a dataset of single-turn question-intent pairs $\mathcal{D}={(x_{i},y_{i})}$ can be easily accessible, where $x_{i}$ denotes the question and $y_{i}\in\mathcal{I}$ is its labelled intent. This dataset is described as single-turn because all $x_{i}$ on their own are standalone complete utterances and are semantically clear for intent inference.
@@ 2.2. Multi-turn
Figure 3. Overview of the Chain-of-Intent framework: using LLM-enhanced HMM to sample intent sequences and generate intent-aware multi-turn dialogues.
@@ 3. Intent-aware Dialogue Generation
Figure 3 shows the detailed workflow of intent-aware dialogue generation, which comprises two main stages: domain knowledge extraction and question generation.
@@ 3.1. Domain Knowledge Extraction
As LLMs do not possess domain-specific knowledge out-of-the-box, this step is imperative to extract the knowledge to be injected into the conversation generation process, bridging the gap for business usage. There are two types of domain-specific knowledge to account for when generating a conversation, namely the number of turns in each conversation and the intent transition from intents of historical questions. These two combined will guide the generation of each conversation.
@@ 3.1.2. Intent Transition Matrix
Other than the number of turns, we also need to know the sensible sequence of intents that users might have in a conversation. Just like the number of turns, the intent transition information is also gathered from the chat logs using Laplace smoothing with $\alpha=0.1$ to handle unseen transitions. One limitation here is that the intent from chat logs is inferred by an intent recognition model trained on the single-turn data $\mathcal{D}$. Each $q_{t}$ is inferred separately without considering $h_{t}$ to obtain $I_{t}$ for the $t$-th turn. That said, by the Law of Large Numbers, if we gather the statistics from a reasonably large number of chat logs, the intent transition matrix obtained will provide a good estimation of true intents. Hereby, the initial intent distribution $P(I_{1})$ is specified in a row matrix, while the intent transition $P(I_{t}|I_{t-1})$ is defined in another $|\mathcal{I}|\times|\mathcal{I}|$ matrix. Note that the $P(I_{t}|I_{t-1})$ here is time invariant, the probability of $P(I_{j}|I_{i})$ will be the same regardless of $t$.
@@ 3.2.1. Intent Sampling with HMM
Using these distributions, we first sample the number of turns $T$ from $P(T)$, then sample the initial intent $I_{1}$ from $P_{\text{init}}$, and subsequent intents $I_{t}$ from $P_{\text{trans}}(\cdot|I_{t-1})$ conditioned on the previous intent.
@@
(1) $P(I_{1:T},q_{1:T})=P(I_{1})P(q_{1}|I_{1})\prod_{t=2}^{T}P(I_{t}|I_{t-1})P(q_{t}|I_{t})$
where $I_{t}$ corresponds to the intent at turn $t$, and $q_{t}$ corresponds to the user’s utterance (question) at turn $t$, with normalization constraints $\sum P(T)=1$ and $\sum P(I_{t}|I_{t-1})=1$. The observation model $P(q_{t}|I_{t})$ is modeled by uniformly sampling a question $X_{t}$ from the set of questions in the single-turn dataset $\mathcal{D}$ that are labeled with intent $I_{t}$:
@@ 3.2.3. Enhancing HMMs with LLMs
To overcome these limitations, we integrate LLMs into the HMM framework, allowing us to generate dialogues that are coherent and context-aware. Specifically, we leverage the LLM’s in-context learning ability to generate each user utterance $q_{t}$ and the corresponding chatbot response $a_{t}$, conditioning on the conversation history up to turn t1t-1.
@@
To incorporate contextual dependencies, we extend the formulation in Equation 2 by conditioning question generation on conversation history:
@@
$\times\prod_{t=2}^{T}P(I_{t}|I_{t-1})P(X_{t}|I_{t})P(a_{t-1}|q_{1:t-1})$
(3) $\times P(q_{t}|I_{t},X_{t},q_{1:t-1},a_{1:t-1})$
@@ 3.3. MINT-E
As ID is the main market with the most traffic online, we observe the top 15 intents based on the question-level ratio in ID generated dataset. The top user intent is to check delivery status, which makes up to 13.8% of the questions. We also observed a long-tailed distribution with the top 10 intents taking up more than 50% of the questions.
@@
Table 1. Comparison of dataset statistics for UDC, MSDialog, and MINT-E.
@@ 4.3.3. Combined Training Objective
Here, $\lambda$ is a hyperparameter that balances the contribution of the contrastive loss. Based on validation performance across three values $\{0.1,0.3,0.5\}$, we set $\lambda=0.3$.
@@ 5.2. Settings for Dialogue Quality Evaluation
We evaluated the quality of the synthetic dialogues using the GPT-4 model. The assessment focused on language fluency, logical flow of topics, and the coherence of questions within each session. Conversations were rated on a scale from 1 to 10, where 1 indicated poor quality and 10 represented excellent quality. The overall score was calculated by averaging the ratings across all evaluated conversations. While this approach may exhibit preference bias toward GPT-generated content, it provides consistent comparative assessment across methods. The prompt for this evaluation can be found in Appendix A.4.
@@ 5.4.1. Quality of Generated Dialogues
Table 3 shows the quality ratings for the generated dialogues across different markets. Our generated dataset, MINT-E, consistently outperformed the baselines in every market. Interestingly, our generated dialogues sometimes received higher average ratings than the golden set, as users often provide brief or out-of-context queries, whereas the language model generates more detailed and contextually relevant questions.
@@
Table 3. Quality ratings of generated dialogues across different methods and markets. Best performance is highlighted in bold.
@@
Table 4. Impact of different training configurations on Multi-Turn Intent Classification accuracy. Best performance (averages across all markets and English-only markets) is highlighted in bold.
@@ 5.4.2. Downstream Task Performance
The impact of different training configurations on the MTIC task is presented in Table 4. Incorporating multi-turn data generally improved performance compared to using only single-turn data. The multi-task learning strategies improved average accuracy by 0.5% on average across 5 markets, particularly in English-speaking markets.
@@ 5.4.3. Discussion and Limitations
Our experiments show that integrating domain knowledge with LLM, as done in MINT-CL, generally enhances dialogue quality and improves performance on downstream tasks. The results suggest that leveraging multi-turn data and multi-task learning is beneficial, especially in markets where the language model performs well.
@@
Table 5. Single-turn (ST) and multi-turn (MT) results on low-resource languages.
@@
Table 6. Comparison of MINT-E statistics with the golden set sampled from real online chat logs.
@@ 7. Related Work
Multi-turn Dialogue Generation (MTDG) emerges from both academia and industry due to its relevance to real-world applications, such as chatbots and information-seeking systems, in production (Liu et al., 2024a; Liu and Fu, 2024). However, they face challenges such as the data sparsity (Zhang et al., 2018c, b; Li et al., 2017). Moreover, MTDG requires more complex information and constraints (Zhang et al., 2020; Chen et al., 2018; Zhang et al., 2018a; Mu and Lim, 2023), posing additional challenges. In general, dialogue generation is categorised into open-domain generation and task-oriented generation (Lv et al., 2023).
@@
Open-domain generation is a context-sensitive process that spans multiple turns, with the model learning to generate appropriate yet open-ended responses based on previous utterances (i.e., the context), which can be categorized into preLLM and postLLM (Soudani et al., 2023). PreLLM methods include the HRED (Serban et al., 2015) series of models. This hierarchical encoder-decoder framework effectively captures contextual information and incorporates latent variables as intermediate states to generate diverse responses, such as variable HRED (Serban et al., 2016b), MrRNN (Serban et al., 2016a) and DialoGPS(Lv et al., 2023). Other related research adopts deep-learning models (Santhanam and Shaikh, 2019) enhanced by internal knowledge to ensure on-topic response generation(Yu et al., 2022; Madotto et al., 2018; Wang et al., 2020a; Mu and Lim, 2024, 2022). Popular architectures include recurrent neural networks (RNNs) (Li et al., 2015), memory networks (Chen et al., 2018) and transformers (Lin et al., 2020; Olabiyi et al., 2023). Those works introduce topic words (Xing et al., 2017), emotion (Song et al., 2019; Zhou et al., 2018), and persona (Zhang, 2018; Zhou et al., 2018) into the generation process to help deliver better dialogue responses. Most of the research still adopts supervised models, which require much training data for better performance. The advent of LLM has created new possibilities for automated text generation. PostLLM methods leverage powerful LLM to utilize prompt or instruction to generate instruction-response pairs automatically. Due to the excellent generation ability of LLM, most research on generation tasks focuses on developing high-quality instruction for multi-turn generation (Wang et al., 2022; Ding et al., 2023) and enhancing the multi-turn instruction following ability with human feedback (Sun et al., 2023; Ouyang et al., 2022). However, these open-domain generation methods do not explicitly adopt domain-specific intents into dialogue generation, so their work is hardly used in other downstream NLP tasks, such as intent or emotion classification tasks.
@@
Task-oriented generation is goal-driven and limited to specific domains for task-oriented dialogue (TOD) (Soudani et al., 2023). It has been widely deployed in many industries as it delivers more efficient customer support (Moghe et al., 2022; Deng et al., 2022). The popular approach centres around user simulation (Lin et al., 2021; Tseng et al., 2021), where two agents interact to process the dialogue history and create a belief state for each side of the conversation, called self-play. They adopt HMM (Cuayáhuitl et al., 2005), partially observable Markov decision processes (POMDP), as a user model for spoken dialogue generation. Recent work also adopts LLM as a user simulator (Kong et al., 2024) with reinforcement learning (Min et al., 2023) and in-context learning framework (Terragni et al., 2023; Askari et al., 2024). This work considers the limited number of dialogue acts, which can’t include all types of intents within the chatbot.
@@
Multi-Turn Intent Classification focuses on understanding user intents within the context of ongoing dialogues. Larson et al. (2019) highlighted the challenge of distinguishing out-of-scope queries, emphasizing the need for robust intent classification models. To enhance interpretability in intent identification, Yang et al. (2020) utilized reinforcement learning over knowledge graphs, demonstrating improved performance in multi-turn settings. Tran et al. (2022) proposed an adaptive global-local context fusion mechanism, effectively integrating dialogue context for enhanced intent classification and slot filling. To address the complexity of multi-turn conversational contexts in the e-commerce domain, Liu et al. (2024b) developed LARA, a Linguistic-Adaptive Retrieval-Augmentation framework across multiple languages. Furthermore, Fan et al. (2022) developed a system leveraging pre-trained language models for fine-grained query understanding, aligning with our focus on intent-aware dialogues in specific domains. Additionally, Hao et al. (2023) introduced a method employing reinforcement learning to construct an intent graph for identification, highlighting the importance of capturing intent transitions in dialogues.